OpenStudy (anonymous):

Solve the equation. Check your answer. 3x – 7 = 5x + 19

OpenStudy (anonymous):

-2x=19+7 -2x=26 x=-26/2 x=-13

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

3x – 7 = 5x + 19 -7 = 5x - 3x + 19 -7 = 2x + 19